OpenStudy (anonymous):

Help plz, Set up an equation to solve the following problem, and then solve it. On a sales trip, Irwin drove 330 miles averaging a certain speed. The return trip was at an average speed that is 11mph faster. Total time for the round trip was 11 hours. Find Irwin’s average speed on each part of the trip.

OpenStudy (anonymous):

well, its not hard , what dont we know , what we do we need to find , assign them to variables

OpenStudy (radar):

Use distance equal rate times time.

OpenStudy (anonymous):

let x equal the "certain speed"

OpenStudy (anonymous):

so assume 330 miles for one way , that means in total he drove 660miles

OpenStudy (anonymous):

we know the total time was 11 but time = distance / speed time for first section = 330 / x times for second section = 330/ ( x+11) so ( 330/x) + [330 / (x+11) = 11

OpenStudy (anonymous):

\[\frac{330}{x} + \frac{330}{x+11} =11\]

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

i c. Thanks

OpenStudy (anonymous):

it has two solutions x=-6, x=55 but because we are dealling with real world ( and yes, it says "speed", not velocity ) so a negative value is not allowed

OpenStudy (anonymous):

so x=55 means for first half average was 55mph, and for second half average was 66mph ( ie 55 +11 )

OpenStudy (radar):

Actually the average speed for the entire trip was 660/11 or 60 mi. So wouldn't that made the avg speed for the first part 54.5 mph and the second part 65.5?

OpenStudy (anonymous):

dnt think so

OpenStudy (radar):

The avg of 54.5 + 65.5 is 120/2=60

OpenStudy (anonymous):

firstly where did u magically get 54.5 and 65.5 from

OpenStudy (radar):

x + (x+11) 660 ---------- = ----- = 60 2 11

OpenStudy (radar):

Nothing magic about it.

OpenStudy (anonymous):

yes.... and.....?

OpenStudy (anonymous):

yes there is , there are alot of different ways to get two numbers to sum to 120 , why did you pick those two numbers? , theres no calculations to suggest why you chose thoise two

OpenStudy (anonymous):

ohh I think i see what you are doing now

OpenStudy (anonymous):

bit of a dangerous way to think

OpenStudy (radar):

x= first avg speed (certain speed) x+11 is the second avg speed (return trip Their sum divided by 2 equals the total avg speed 660/11 is also the total avg speed. Take it from there

OpenStudy (radar):

Just sayin your way seems correct just wondering why they dont come out the same. Did you round off or take a root of or whatever/

OpenStudy (anonymous):

no, the reason why they dont come out correct is because your way of thinking is wrong :P

OpenStudy (radar):

Well they came out close.lol gotta run

OpenStudy (anonymous):

:\, the average of a sum is not the sum of the avergaes

OpenStudy (radar):

Gotcha, I see the error of my logic.